A random variable X is normally distributed with a mean of 45 and standard deviation of 3.If the area to the right of Z is 0.9292, what will be the value of its random variable X


1
Expert's answer
2022-04-14T14:00:19-0400

"\mu=45,\sigma=3,P(Z>z)=0.9292,\\
P(Z<z)=1-0.9292=0.0708."

From z-table we find "z=-1.47."

So, "P(Z<z)=P(X<x), \\
z=\cfrac{x-\mu} {\sigma},\\
x=z\sigma+\mu=-1.47\cdot3+45=40.59."